OpenStudy (anonymous):

Simplify radical : square root of 363

OpenStudy (kropot72):

\[\sqrt{363}=\sqrt{3\times 11\times 11}=\ ?\] You can take 11 * 11 outside the radical sign, but don't forget to take the square root when you do.

OpenStudy (anonymous):

6 square root 5

OpenStudy (kropot72):

\[\sqrt{363}=11\sqrt{3}\]